the body temperature of healthy children is normally distributed with mean 98.6 F and standard deviation 0.9 F.
If one healthy child is selected at random, what is the probability that the body temperature is more than 99.6 F?
If six healthy children are chosen at random, what is the probability that the mean body temperature x for all six is more than 99.6F?

Answer by ewatrrr(24785)

 
Hi,
body temperature of healthy children is normally distributed with mean 98.6 F and standard deviation 0.9 F.
If one healthy child is selected at random:
P(temp>99.6)= P(z> (99.9-98.6)/.9) = P(z> 1.7778) = .0377
If six healthy children are chosen at random:
P(6 with temp >99.6) = (.0377)^6